Use the Clausius-Clapeyron equation to calculate the vapor pressure for the following samples. Assume the amount of each sample is 1 mole.
a. Water at 63 C.
b. Ethanol at 23 C.
c. Chloroform at 52 °C.
1
Expert's answer
2019-10-21T07:16:25-0400

The vapor pressure of any liquid at its boiling point is 1 atm.

(a) for water, "P_1=1 atm,T_1=373 \ K,T_2=63\degree C=63+273=333\ K"

"\triangle H_{vap} =40.65KJ/mol"

By using Clausius-Clayperon equation,

"ln\frac{P_2}{P_1}=\frac{\triangle H_{vap}}{R}(\frac{1}{T_1}-\frac{1}{T_2})"

"ln\frac{P_2}{1}=\frac{40.65\times 10^3}{8.314}(\frac{1}{373}-\frac{1}{333})"

"P_2=0.2071 \ atm"

(b) for ethanol, "P_1=1atm,T_1(B.P.)=78.37 \degree C=351.37 K," "T_2=23\degree C=296 K"

"\triangle H_{vap}=38.6 KJ/mol"

By using Clausius-Clayperon equation,

"ln\frac{P_2}{P_1}=\frac{\triangle H_{vap}}{R}(\frac{1}{T_1}-\frac{1}{T_2})"

"ln\frac{P_2}{1}=\frac{38.6\times 10^3}{8.314}(\frac{1}{351.37}-\frac{1}{296})"

"P_2=0.84441 \ atm"

(c) For chloroform,"P_1=1atm,T_1(B.P.)=61.2 \degree C=334.2 K," "T_2=52\degree C=325 K"

"\triangle H_{vap}=31.4 KJ/mol"

By using Clausius-Clayperon equation,

"ln\frac{P_2}{P_1}=\frac{\triangle H_{vap}}{R}(\frac{1}{T_1}-\frac{1}{T_2})"

"ln\frac{P_2}{1}=\frac{31.4\times 10^3}{8.314}(\frac{1}{334.2}-\frac{1}{325})"

"P_2=0.72622 \ atm"